OpenStudy (anonymous):

Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x2 − 4x y = −2x2 + 8x − 12

OpenStudy (anonymous):

y = x² − 4x ⇒ opens upward = x² − 4x + 4 − 4 = (x − 2)² − 4 ⇒ vertex is at (2, −4) y = −2x² + 8x − 12 ⇒ opens downward, narrower than first one because x²-coefficient is farther away from 0 y = −2(x² − 4x + 6) y = −2(x² − 4x + 4 + 2) y = −2(x − 2)² − 4 ⇒ vertex is at (2, −4), same place as first one ...once you have this information, this shouldn't be a problem.